# Learning Regularization for Graph Inverse Problems

### v1

@@ Abstract
In recent years, Graph Neural Networks (GNNs) have been utilized for various applications ranging from drug discovery to network design and social networks. In many applications, it is impossible to observe some properties of the graph directly; instead, noisy and indirect measurements of these properties are available. These scenarios are coined as Graph Inverse Problems (GRIP). In this work, we introduce a framework leveraging GNNs to solve GRIPs. The framework is based on a combination of likelihood and prior terms, which are used to find a solution that fits the data while adhering to learned prior information. Specifically, we propose to combine recent deep learning techniques that were developed for inverse problems, together with GNN architectures, to formulate and solve GRIP. We study our approach on a number of representative problems that demonstrate the effectiveness of the framework.
@@ 1 Introduction
In recent years, graph machine learning frameworks were developed, predominantly, Graph Neural Networks (Scarselli et al. 2008; Bronstein et al. 2021), which are able to model and learn complex patterns in graph data. These recent advancements make it possible to perform a wide array of applications, such as node classification (Kipf and Welling 2016; Defferrard, Bresson, and Vandergheynst 2016), community detection (Chen, Li, and Bruna 2019), drug discovery (Jiang et al. 2021), and solving combinatorial problems (Schuetz, Brubaker, and Katzgraber 2022).
@@
To address this question and its challenges, we begin by defining Graph Inverse Problems (GRIP), highlighting a key difference between classical inverse problems that are commonly solved on structured grids and those on graphs. This difference directly stems from the presence of the graph, which provides additional structure. Furthermore, in many GRIP cases, there exist additional features on the graphs (throughout this paper, we will refer to these features as meta-data) which may not be directly related to the observed data associated with the inverse problem, but can be leveraged to obtain better estimates for the solution of the inverse problem. Following the definition of such problems, and a number of examples, we discuss several viable approaches, from classical to neural methods, and evaluate them on several datasets and GRIPs.
@@
Main Contributions.
This paper advances the bridge between the field of inverse problems and graph machine learning by an overarching methodology for the solution of inverse problems that reside on graphs, that is, GRIP. By integrating the principles of learned regularization techniques (Adler and Öktem 2017; Mardani et al. 2018), with the flexibility of GNNs (Khemani et al. 2024), this work aims to develop a unified framework that leverages the strengths of both approaches, to solve GRIP. We demonstrate our framework on several key GRIPs, such as graph feature completion, source estimation, inverse graph transport, and the nonlinear edge recovery problem. We note that the proposed framework not only enhances the existing techniques for GRIP but also broadens the scope of GNN applications.
@@ 2 Graph Inverse Problems
where $F:{\cal X}\rightarrow{\cal D}$ is the forward map of the problem that maps the states ${\bf x}$ that reside on the graph to the data space ${\cal D}$. The vector ${\boldsymbol{\varepsilon}}$ is some noise that, for simplicity, is assumed to be $\sigma\sim N(0,\sigma^{2}{\bf I})$. While for some problems the noise can be substantial, for many practical problems (e.g. graph segmentation) the noise can be negligent.
@@ 2.3 Inverse Graph Transport (Figure 3)
Remark.
The problems defined in Equations (2)–(4) are linear. Next, we show an example of a nonlinear problem, which is considered more complex (Manríquez et al. 2021).
@@ 2.4 Edge Property Recovery (Figure 4)
Consider the case of source estimation as defined in Equation (3), but with known source and unknown edge properties where
@@
For the forward problems presented in this section, their inverse problems are assumed to be ill-posed. That is, the operator $F$ in Equation (1) has a large, non-trivial null space or, such that the singular values of its Jacobian quickly decay to 00. This implies that a small perturbation in the data can result in a large perturbation in the estimated solution (see, e.g., (Hansen 1997)). Therefore, to obtain an estimate of the solution, some regularization is required. Such regularization can vary from classical techniques like Tikhonov regularization (Tikhonov 1950) (similar to weight-decay in neural networks), smoothness-based regularization (Nadler, Srebro, and Zhou 2009), total-variation (Rudin, Osher, and Fatemi 1992), as well as learnable neural regularizers (Adler and Öktem 2017). In previous works on graph inverse problem (Huang et al. 2023), such a regularization was achieved implicitly by learning a map, $F^{{\dagger}}$, from the data ${\bf d}^{\rm obs}$ to the solution ${\bf x}$. Nonetheless, as has been shown for other, non-graph inverse problems (Adler and Öktem 2017; Eliasof, Haber, and Treister 2023), even when the data fit is a part of the training, at inference, such a map can lead to reconstructions that do not fit the observed data, – which can lead to non-feasible solutions. Mathematically, these findings mean that the residual between the predicted solution and the observed data, defined as:
@@ 3 Methodologies for Inverse Problems
Goal.
The objective in solving Inverse Problems is to estimate a solution $\widehat{\bf x}$ that (i) fits the data in Equation (1) to some prescribed error, and (ii) adheres to a priori information given by a regularizer. Two approaches to solving such problems are commonly used. The first, is based on a variational method and the second is to embed the solution with an (inverse) scale-space dynamical system. We now shortly review both techniques and their behavior.
Variational Techniques.
Variational methods formulate inverse problems as an optimization task, seeking to minimize a functional that balances fidelity to the observed data with the imposition of prior knowledge or regularization (see, e.g., (Engl, Hanke, and Neubauer 1996; Calvetti and Somersalo 2005; Tenorio et al. 2011)). They have a strong relation to the Maximum Aposteriori Estimate (MAP) when considering a Bayesian approach (Tarantola 1987). In the context of inverse problems, the objective functional includes a data fitting term, which ensures that the solution is consistent with the observed data, and a regularization term, which incorporates prior assumptions about the solution, such as smoothness, sparsity, or other structural properties. Here, we consider regularization techniques where the estimated $\widehat{\bf x}$ is obtained by solving a regularized data fitting problem:
@@
Inverse Scale Space Methods.
One of the difficulties in solving the optimization proposed in Variational Methods, as in Equation (6a), is the appropriate choice of the relative weight $\alpha$, between the data fitting and regularization terms (Nagy and Hansen 2006). Inverse Scale space methods (Burger et al. 2006) propose an alternative approach by casting the optimization problem to a dynamical system that starts with a coarse approximation of the solution and progressively incorporates finer details. By initially focusing on the low-frequency details and gradually refining the solution to encode high-frequency details, inverse scale space techniques can achieve a balance between fidelity to the data and the incorporation of prior knowledge. The choice of the regularization parameter is replaced by the choice of the stopping time, which is often more straightforward.
@@
${\frac{d{\bf z}}{dt}}={\bf E}(t)^{\top}{\bf J}(t)^{\top}\left({\bf d}^{\rm obs}-F({\bf E}(t){\bf z})\right)-s({\bf z},{\boldsymbol{\theta}}(t)),$
${\bf z}(0)={\bf z}_{0},$ (7)
from time 00 to time $T$, which is determined by the level of fitting the data. Here, ${\bf E}(t)$ is a time-dependent embedding, ${\bf J}(t)$ is the Jacobian of $F$ with respect to its argument, and $s$ is a function referred to as the score. That is, in order to use a scale space approach for the solution of the problem, one is required to choose time-dependent embedding and a score function, $s$. The flexibility of a time-dependent embedding ${\bf E}$ and regularization $s$ was shown to overcome issues such as local minima (Eliasof, Haber, and Treister 2024).
We note that, while Inverse scale space methods offer more flexibility, they have less theoretical understanding compared with variational methods. Additionally, scale space methods can be reduced to solving the optimization problem in Equation (6a) using gradient descent, if we choose $s=\nabla{\bf R}$ and ${\bf E}$ to be time-invariant; however, typically, in Scale Space methods, $s$ and ${\bf E}$ are time-dependent.
@@ 4 A Framework for Solving GRIP
(i) Graph-Informed Regularizers.
Because the problem is defined on a graph, it is possible and desirable to leverage the graph structure to obtain additional insights on the support of the problem. Regularization techniques can incorporate the connectivity and topology of the graph, enabling more informed and precise recovery of the property ${\bf x}$. For example, smoothness constraints can be imposed, effectively promoting solutions where neighboring nodes have similar values, which is particularly useful in applications like homophilic node classification (Zhou et al. 2004).
(ii) Utilization of Meta-Data in Graphs.
While the observed data ${\bf d}^{\rm obs}$, defined in Equation (1), depends solely on states ${\bf x}$, in many cases, each node or edge in the graph is associated with additional meta-data ${\bf f}_{\rm{M}}$. This meta-data, which may include attributes such as node positions, edge weights, or additional features, is not required to compute ${\bf d}^{\rm obs}$, but it provides valuable supplementary information that can enhance the regularization process. For example, in applications related to 3D Computer Vision, where point clouds and graphs are processed, it is common to have multiple types of features, such as $xyz$ coordinates, point normals, node-wise labels, as well as a global label of the graph. Additionally, we note that in the absence of observed data ${\bf d}^{\rm{obs}}$, the proposed framework reduces to standard GNN architectures, as discussed below. Overall, this two-fold consideration of the graph structure and meta-data enables a more holistic approach to solving GRIP, as we now discuss.
@@ 4.1 Classical Graph Regularization
Laplacian Regularization.
A classical regularizer that has been used extensively, especially in the case of graph completion (Nadler, Srebro, and Zhou 2009) is based on the graph Laplacian ${\bf L}={\bf D}-{\bf A}$. Given the graph Laplacian ${\bf L}$, the Laplacian smoothness regularizer is defined as
@@
Tikhonov Regularization.
The Tikhonov Regularization technique (Tikhonov and Arsenin 1977), which promotes solutions with low norms, is defined by replacing the Laplacian ${\bf L}$ in Equation (8) by the identity matrix ${\bf I}$.
@@ 4.2 Neural Graph Regularization
Variational Methods.
We consider the use of variational regularization, discussed in Section 3, where the key idea is to learn the regularization term using a GNN. In particular, as previously discussed, when working with graph data, there often exist additional, meta-data, ${\bf f}_{M}$ for each node. Thus, we modify a recent regularization operator, proposed in (Eliasof, Haber, and Treister 2023) to include meta-data, such that this regularization reads:
${{\bf R}}({\bf x}^{(1)},\ldots,{\bf x}^{(L)},{\bf f}_{M},\{{\boldsymbol{\theta}}{{}_{l}}\}_{l=1}^{L})$
$=\underbrace{\sum_{l=1}^{L-1}{\frac{1}{2}}\|{\bf x}^{(l+1)}-{\bf x}^{(l)}\|^{2}}_{Kinetic}\ +\ \underbrace{\sum_{l=1}^{L-1}{\boldsymbol{\Phi}}({\bf x}^{(l)},{\bf f}_{M},{\cal G},{\boldsymbol{\theta}}_{l})}_{Potential}.$
Here, the solution ${\bf x}={\bf x}^{(L)}$ is embedded in the hidden states ${\bf x}^{(1)},\ldots,{\bf x}^{(L)}$ and $\hat{{\bf x}}={\bf x}^{(L)}$ is the predicted solution.
As shown in Equation (4.2), the regularization employs a kinetic energy term and a potential energy term.
We now show the effect of the meta-data ${\bf f}_{M}$ in this regularizer. To this end, consider the minimization problem in Equation (6a), when there is no observed data ${\bf d}^{\rm{obs}}$. That is, we only minimize the regularization term. Differentiating Equation (4.2) with respect to ${\bf x}$ we obtain:
${\bf x}_{k+1}=2{\bf x}_{k}-{\bf x}_{k-1}+\nabla_{{\bf x}_{i}}{\boldsymbol{\Phi}}({\bf x}_{i},{\bf f}_{M},{\cal G},{\boldsymbol{\theta}}_{i}).$ (12)
@@
Inverse Scale Space.
The second technique we experiment with is a modification of a learned inverse scale space technique proposed in (Eliasof, Haber, and Treister 2024) for solving inverse problems on a structured grid. We adapt it to operate on graph problems. The dynamical system (10) is simply discretized in two steps, obtaining:
@@
${\bf z}_{k+1}={\bf z}_{k+{\frac{1}{2}}}-s({\bf z}_{k+{\frac{1}{2}}},{\bf f}_{M},{\cal G},{\boldsymbol{\theta}}_{k},t_{k})$ (13b)
@@
Proximal Methods.
Another approach for solving inverse problems that blends scale-space and variational methods was proposed in (Adler and Öktem 2017; Mardani et al. 2018). Here, we modify the technique for graph learning.
@@
This approach leads to a two-stage network, where in the first step, a gradient step is applied (Equation (14a)), which aims to fit the data, and in the second step, a proximal step is applied (Equation (4.2)). In Adler and Öktem (2017), the iteration in Equation (4.2) shares the same parameters ${\boldsymbol{\theta}}$ for every proximal step. In Mardani et al. (2018), the iteration is unrolled, to learn a parameter ${\boldsymbol{\theta}}_{k}$ per-iteration. This choice unties the link to variational methods but increases the flexibility of the method. In our experiments we follow Mardani et al. (2018), and modify the method proposed there to operate on graphs. We call this network Prox-GNN.
Mathematically, uniting the two steps in Equation (14) leads to Prox-GNN, as follows:
@@
where $s(\cdot)$ is a GNN that uses both observed data ${\bf d}^{\rm{obs}}$ and the meta-data to solve the problem.
It is interesting to observe that, in the absence of any observed data, this network turns into a standard feed-forward GNN that uses the mata-data to predict ${\bf x}$.

### v2

@@ Abstract
In recent years, Graph Neural Networks (GNNs) have been utilized for various applications ranging from drug discovery to network design and social networks. In many applications, it is impossible to observe some properties of the graph directly; instead, noisy and indirect measurements of these properties are available. These scenarios are coined as Graph Inverse Problems (GRIPs). In this work, we introduce a framework leveraging GNNs to solve GRIPs. The framework is based on a combination of likelihood and prior terms, which are used to find a solution that fits the data while adhering to learned prior information. Specifically, we propose to combine recent deep learning techniques that were developed for inverse problems, together with GNN architectures, to formulate and solve GRIPs. We study our approach on a number of representative problems that demonstrate the effectiveness of the framework.
@@ 1 Introduction
In recent years, graph machine learning frameworks were developed, predominantly, Graph Neural Networks (Scarselli et al. 2008; Bronstein et al. 2021), which are able to model and learn complex patterns in graph data. These recent advancements make it possible to perform a wide array of applications, such as node classification (Kipf and Welling 2016; Defferrard, Bresson, and Vandergheynst 2016), community detection (Chen, Li, and Bruna 2019), drug discovery (Jiang et al. 2021), and solving combinatorial problems (Schuetz, Brubaker, and Katzgraber 2022; Eliasof and Haber 2024).
@@
To address this question and its challenges, we begin by defining Graph Inverse Problems (GRIPs), highlighting a key difference between classical inverse problems that are commonly solved on structured grids and those on graphs. This difference directly stems from the presence of the graph, which provides additional structure. Furthermore, in many GRIP cases, there exist additional features on the graphs (throughout this paper, we will refer to these features as meta-data) which may not be directly related to the observed data associated with the inverse problem, but can be leveraged to obtain better estimates for the solution of the inverse problem. Following the definition of such problems, and a number of examples, we discuss several viable approaches, from classical to neural methods, and evaluate them on several datasets and GRIPs.
@@
Main Contributions. This paper advances the bridge between the field of inverse problems and graph machine learning by an overarching methodology for the solution of inverse problems that reside on graphs, that is, GRIP. By integrating the principles of learned regularization techniques (Adler and Öktem 2017; Mardani et al. 2018), with the flexibility of GNNs (Khemani et al. 2024), this work aims to develop a unified framework that leverages the strengths of both approaches, to solve GRIP. We demonstrate our framework on several key GRIPs, such as graph feature completion, source estimation, inverse graph transport, and the nonlinear edge recovery problem. We note that the proposed framework not only enhances the existing techniques for GRIP but also broadens the scope of GNN applications.
@@ 2 Graph Inverse Problems
where $F:{\cal X}\rightarrow{\cal D}$ is the forward map of the problem that maps the states ${\bf x}$ that reside on the graph to the data space ${\cal D}$. The vector ${\boldsymbol{\varepsilon}}$ is some noise that, for simplicity, is assumed to be $\sigma\sim N(0,\sigma^{2}{\bf I})$. While for some problems, the noise can be substantial, for many practical problems (e.g., graph segmentation), the noise can be negligent.
@@ 2.3 Inverse Graph Transport (Figure 3)
Remark. The problems defined in Equations (2)–(4) are linear. Next, we show an example of a nonlinear problem, which is considered more complex (Manríquez et al. 2021).
@@ 2.4 Edge Property Recovery (Figure 4)
Consider the case of source estimation in Equation (3), but with known source and unknown edge properties where
@@
For the forward problems presented in this section, their inverse problems are assumed to be ill-posed. That is, the operator $F$ in Equation (1) has a large, non-trivial null space or, such that the singular values of its Jacobian quickly decay to 00. This implies that a small perturbation in the data can result in a large perturbation in the estimated solution (see, e.g., (Hansen 1997)). Therefore, to obtain an estimate of the solution, some regularization is required. Such regularization can vary from classical techniques like Tikhonov regularization (Tikhonov 1950) (similar to weight-decay in neural networks), smoothness-based regularization (Nadler, Srebro, and Zhou 2009), total-variation (Rudin, Osher, and Fatemi 1992), as well as learnable neural regularizers (Adler and Öktem 2017). In previous works on graph inverse problem (Huang et al. 2023), such a regularization was achieved implicitly by learning a map, $F^{{\dagger}}$, from the data ${\bf d}^{\rm obs}$ to the solution ${\bf x}$. Nonetheless, as has been shown for other, non-graph inverse problems (Adler and Öktem 2017; Eliasof, Haber, and Treister 2023), even when the data fit is a part of the training, at inference, such a map can lead to reconstructions that do not fit the observed data – which can lead to non-feasible solutions. Mathematically, these findings mean that the residual between the predicted solution and the observed data, defined as:
@@ 3 Methodologies for Inverse Problems
Goal. The objective in solving Inverse Problems is to estimate a solution $\widehat{\bf x}$ that (i) fits the data in Equation (1) to some prescribed error, and (ii) adheres to a-priori information given by a regularizer. Two approaches to solving such problems are commonly used. The first, is based on a variational method and the second is to embed the solution with an (inverse) scale-space dynamical system. We now shortly review both techniques and their behavior.
Variational Techniques. Variational methods formulate inverse problems as an optimization task, seeking to minimize a functional that balances fidelity to the observed data with the imposition of prior knowledge or regularization (see, e.g., (Engl, Hanke, and Neubauer 1996; Calvetti and Somersalo 2005; Tenorio et al. 2011)). They have a strong relation to the Maximum Aposteriori Estimate (MAP) when considering a Bayesian approach (Tarantola 1987). In the context of inverse problems, the objective functional includes a data fitting term, which ensures that the solution is consistent with the observed data, and a regularization term, which incorporates prior assumptions about the solution, such as smoothness, sparsity, or other structural properties. Here, we consider regularization techniques where the estimated $\widehat{\bf x}$ is obtained by solving a regularized data fitting problem:
@@
Inverse Scale-Space Methods. One of the difficulties in solving the optimization proposed in Variational Methods, as in Equation (6a), is the appropriate choice of the relative weight $\alpha$, between the data fitting and regularization terms (Nagy and Hansen 2006). Inverse scale-space methods (Burger et al. 2006) propose an alternative approach by casting the optimization problem to a dynamical system that starts with a coarse approximation of the solution and progressively incorporates finer details. By initially focusing on the low-frequency details and gradually refining the solution to encode high-frequency details, inverse scale-space techniques can achieve a balance between fidelity to the data and the incorporation of prior knowledge. The choice of the regularization parameter is replaced by the choice of the stopping time, which is often more straightforward.
@@
${\frac{d{\bf z}}{dt}}={\bf E}(t)^{\top}{\bf J}(t)^{\top}\left({\bf d}^{\rm obs}-F({\bf E}(t){\bf z})\right)-s({\bf z},{\boldsymbol{\theta}}(t)),$ (7)
from time 00 to time $T$, with ${\bf z}(0)={\bf z}_{0}$, which is determined by the level of fitting the data. Here, ${\bf E}(t)$ is a time-dependent embedding, ${\bf J}(t)$ is the Jacobian of $F$ with respect to its argument, and $s$ is a function referred to as the score. That is, in order to use a scale space approach for the solution of the problem, one is required to choose time-dependent embedding and a score function, $s$. The flexibility of a time-dependent embedding ${\bf E}$ and regularization $s$ was shown to overcome issues such as local minima (Eliasof, Haber, and Treister 2024). We note that, while Inverse scale-space methods offer more flexibility, they have less theoretical understanding compared with variational methods. Additionally, scale space methods can be reduced to solving the problem in Equation (6a) using gradient descent, if we choose $s=\nabla{\bf R}$ and ${\bf E}$ to be time-invariant; however, typically, in Scale Space methods, $s$ and ${\bf E}$ are time-dependent.
@@ 4 A Framework for Solving GRIP
(i) Graph-Informed Regularizers. Because the problem is defined on a graph, it is possible and desirable to leverage the graph structure to obtain additional insights on the support of the problem. Regularization techniques can incorporate the connectivity and topology of the graph, enabling more informed and precise recovery of the property ${\bf x}$. For example, smoothness constraints can be imposed, effectively promoting solutions where neighboring nodes have similar values, which is particularly useful in applications like homophilic node classification (Zhou et al. 2004).
(ii) Utilization of Meta-Data in Graphs. While the observed data ${\bf d}^{\rm obs}$, defined in Equation (1), depends solely on states ${\bf x}$, in many cases, each node or edge in the graph is associated with additional meta-data ${\bf f}_{\rm{M}}$. This meta-data, which may include attributes such as node positions, edge weights, or additional features, is not required to compute ${\bf d}^{\rm obs}$, but it provides valuable supplementary information that can enhance the regularization process. For example, in applications related to 3D Computer Vision, where point clouds and graphs are processed, it is common to have multiple types of features, such as $xyz$ coordinates, point normals, node-wise labels, as well as a global label of the graph. Additionally, we note that in the absence of observed data ${\bf d}^{\rm{obs}}$, the proposed framework reduces to standard GNN architectures, as discussed below. Overall, this two-fold consideration of the graph structure and meta-data enables a more holistic approach to solving GRIP, as we now discuss.
@@ 4.1 Classical Graph Regularization
Laplacian Regularization. A classical regularizer that has been used extensively, especially in the case of graph completion (Nadler, Srebro, and Zhou 2009) is based on the graph Laplacian ${\bf L}={\bf D}-{\bf A}$. Given the graph Laplacian ${\bf L}$, the Laplacian smoothness regularizer is defined as
@@
Tikhonov Regularization. The Tikhonov Regularization technique (Tikhonov and Arsenin 1977), which promotes solutions with low norms, is defined by replacing the Laplacian ${\bf L}$ in Equation (8) by the identity matrix ${\bf I}$.
@@ 4.2 Neural Graph Regularization
Variational Methods. We consider the use of variational regularization, discussed in Section 3, where the key idea is to learn the regularization term using a GNN. In particular, as previously discussed, when working with graph data, there often exist additional, meta-data, ${\bf f}_{M}$ for each node. Thus, we modify a recent regularization operator, proposed in (Eliasof, Haber, and Treister 2023) to include meta-data, such that this regularization reads:
${{\bf R}}({\bf x}_{1},\ldots,{\bf x}_{L},{\bf f}_{M},\{{\boldsymbol{\theta}}{{}_{l}}\}_{l=1}^{L})$ (11)
$=\underbrace{\sum_{l=1}^{L-1}{\frac{1}{2}}\|{\bf x}_{l+1}-{\bf x}_{l}\|^{2}}_{Kinetic}\ +\ \underbrace{\sum_{l=1}^{L-1}{\boldsymbol{\Phi}}({\bf x}_{l},{\bf f}_{M},{\cal G},{\boldsymbol{\theta}}_{l})}_{Potential}.$
Here, the solution ${\bf x}={\bf x}_{L}$ is embedded in the hidden states ${\bf x}_{1},\ldots,{\bf x}_{L}$ and $\hat{{\bf x}}={\bf x}_{L}$ is the predicted solution.
As shown in Equation (11), the regularization employs a kinetic energy term and a potential energy term.
We now show the effect of the meta-data ${\bf f}_{M}$ in this regularizer. To this end, consider the minimization problem in Equation (6a), when there is no observed data ${\bf d}^{\rm{obs}}$. That is, we only minimize the regularization term. Differentiating Equation (11) with respect to ${\bf x}$ yields the leapfrog scheme:
${\bf x}_{k+1}=2{\bf x}_{k}-{\bf x}_{k-1}+\nabla_{{\bf x}_{k}}{\boldsymbol{\Phi}}({\bf x}_{k},{\bf f}_{M},{\cal G},{\boldsymbol{\theta}}_{k}).$ (12)
@@
Inverse Scale-Space. The second technique we experiment with is a modification of a learned inverse scale-space technique proposed in (Eliasof, Haber, and Treister 2024) for solving inverse problems on a structured grid. We adapt it to operate on graph problems. The dynamical system in Equation (7) is simply discretized in two steps, obtaining:
@@
${\bf z}_{k+1}={\bf z}_{k+{\frac{1}{2}}}-s({\bf z}_{k+{\frac{1}{2}}},{\bf f}_{M},{\cal G},{\boldsymbol{\theta}}_{k},t_{k}).$ (13b)
@@
Proximal Methods. Another approach for solving inverse problems that blends scale-space and variational methods was proposed in (Adler and Öktem 2017; Mardani et al. 2018). Here, we modify the technique for graph learning.
@@
This approach leads to a two-stage network, where in the first step, a gradient step is applied (Equation (14a)), which aims to fit the data, and in the second step, a proximal step is applied (Equation (4.2)). In Adler and Öktem (2017), the iteration in Equation (4.2) shares the same parameters ${\boldsymbol{\theta}}$ for every proximal step. In Mardani et al. (2018), the iteration is unrolled, to learn a parameter ${\boldsymbol{\theta}}_{k}$ per-iteration. This choice unties the link to variational methods but increases the flexibility of the method. In our experiments we follow Mardani et al. (2018), and modify the method proposed there to operate on graphs. We call this network Prox-GNN, which unites the two steps in Equation (14) as follows:
@@
where $s(\cdot)$ is a GNN that uses observed data ${\bf d}^{\rm{obs}}$ and the meta-data to solve the problem. It is interesting to observe that, in the absence of observed data, we obtain a standard feed-forward GNN that uses the meta-data to predict ${\bf x}$.